OpenStudy (anonymous):

this has me confused, please help solve.. a bag contains a total of 9 batteries, of which 5 are defective selecting 2 at random without replacement determine the probability that none of the batteries you select are good

OpenStudy (anonymous):

combination problem

OpenStudy (anonymous):

so would the answer be 5 /18

OpenStudy (anonymous):

well, almost :)

OpenStudy (anonymous):

can you evaluate this: 5C2/9C2

OpenStudy (anonymous):

? ok where did i go wrong?

OpenStudy (anonymous):

not completely.

OpenStudy (anonymous):

none are good means both defective right? so take \[\frac{5}{9}\times \frac{4}{8}\]

OpenStudy (anonymous):

probability first one is defective is 5/9, probability second one is defective GIVEN that the first one was is 4/8 (clear yes? you have 8 batteries left of which 4 are defective) AND means multiply them together

OpenStudy (anonymous):

in fact i too get \[\frac{5}{18}\] so unless i made a mistake (possible for sure) you are correct

OpenStudy (anonymous):

i am sorry, @gensono2, i am a victum once again of my laziness, i had the answer as being 10/36, which means i did not simplify it 5/18.

OpenStudy (anonymous):

you were correct

OpenStudy (anonymous):

thank you. i do appreciate your help.. math and me just sometimes just dont get along (smiling)